# Achieving Collective Welfare in Multi-Agent Reinforcement Learning via Suggestion Sharing

### v1

@@ 6 Discussion and Conclusion
In this work, we addressed the challenge of achieving collective welfare in scenarios where individual interests may conflict with collective objectives. We introduced a MARL method based on Suggestion Sharing (SS), designed for situations where agents lack access to others’ rewards and policies, and traditional methods relying on sharing rewards, values, or policy models are infeasible. SS enables agents to incorporate their individual interests into action suggestions for other agents, facilitating implicit inferences about collective interests and learning policies that promote collective welfare.
@@ Acknowledgments
GM acknowledges support from a UKRI AI Turing Acceleration Fellowship (EPSRC EP/V024868/1).
@@ A.1.3 Proof of Theorem 1
A.2 Algorithm Pseudocode and Illustration
Algorithm 1 MARL with Suggestion Sharing (SS)
  for episode = 1 to $E$ do
     $\mathcal{D}_{i}\leftarrow\phi,\forall i$
     Observe initial state $s_{1}$
     for $t=1$ to $T$ do
        Execute action $a_{t}^{i}\in\pi^{ii}(\cdot|s_{t})$
        Observe reward $r_{t}^{i}$ and next state $s_{t+1}$
        Store $(s_{t},a_{t}^{i},r_{t}^{i},s_{t+1})\in\mathcal{D}_{i}$
     end for
     for iteration = 1 to $K$ do
        for each agent $i$ do
        end for
        for $i=1$ to $N$ do
           Update $\boldsymbol{\tilde{\pi}}^{i}$ using Eq 16
           Update $V_{i}$ using Eq 18
           $\boldsymbol{\tilde{\pi}}^{i}_{old}\leftarrow\boldsymbol{\tilde{\pi}}^{i}$
        end for
     end for
  end for
@@ A.5 Hyperparameters
The hyperparameters used in our experiments are listed in Tables 1 and 2.

### v2

@@ 6 Discussion and Conclusion
In this work, we addressed the challenge of achieving collective welfare in scenarios where individual interests may conflict with collective objectives. We proposed a Suggestion-Sharing-based MARL method, designed for situations where agents lack access to others’ rewards and policies, and traditional methods relying on sharing rewards, values, or policy models are infeasible. SS enables agents to incorporate their individual interests into action suggestions for other agents. Taking into account the suggestions shared by others when learning individual policies can facilitate implicit inferences about collective interests and then facilitate learning policies that can promote collective welfare.
@@ Acknowledgments
Giovanni Montana acknowledges support from a UKRI AI Turing Acceleration Fellowship (EPSRC EP/V024868/1).
@@ A.1.3 Proof of Theorem 1
A.2 Algorithm Illustration
Fig. 6 shows an illustration of our SS-based MARL algorithm.
@@ A.5 Hyperparameters
The hyperparameters used in our experiments are listed in Tables 2 and 3.